News and Updates

News and Updates

About the blog

Read below our regular monthly / quarterly updates of work in progress and operations of the diesel locomotives based at the Bodmin and Wenford Railway.

Read more about the loco's and the group on our full website by clicking here.

Jan - Feb 2017

UpdatesPosted by Administrator Sun, February 26, 2017 13:30:13

Loco Status

47306 ‘The Sapper’ - In overhaul.

D3452 – Stored awaiting overhaul.

P403D ‘Denise’, 08444, 33110, 37142 & 50042 ‘Triumph’ – In Traffic.

Operational dates this period.

None this period.

Various dates – 08444, D3452 & 33110 - Permanent way and station shunt duties.

Future running dates.

18th March x5 return runs 37142 top & tail with steam Red TT.

19th March x5 return runs 37142 top & tail with steam Red TT.

9th April x2 return runs Loco TBD Yellow TT.

23rd April x2 return runs Loco TBD Yellow TT.

29th April x6 return runs Loco TBD Lilac TT.

Maintenance for Service Duties

33110

Used for a couple of minor shunts and 50042 test run see below.

37142

Over Xmas 37142 saw a trip to the engine shed pit for bogie, under chassis, traction motor inspections, cleaning and servicing. Further work has seen an ‘A’ & ‘B’ Exam completed and the two leaking axle box gaskets renewed along with a grease purge ready for the 2017 season start.

With its radiator fan cover in place over the winter, 37142 has stayed relatively dry and fired up Mid-February on a cold day with flakes of snow falling. No issues at all found on a test run to Bodmin Parkway and back.

Photo James Goodhall, 37142 warming up from a 'Cold' start

50042 ‘Triumph’

Triumph was not forgotten over Xmas with a trip down to Parkway shed using 33110, saw us exchange out the old KV10 main generator power control module and fit the newly constructed one.

Whilst two members worked on this, another two carried out parts of an ‘A’ & ‘B’ Exam plus some servicing. Coupling up 33110 to 50042, the latter was extracted from the shed, prepared and started with testing of the new KV10 carried out. Declared working, 50042 was given a work out with 33110 acting as a load, whilst more monitoring of 50042 power handling was undertaken on our 1:40 gradient. The new KV10 replacement works well and early signs are that 50042 is handling power better than it has done since preservation with us at B&WR.

Photo 50042 New KV10 module installed and first checks being carried out on it.

In Overhaul

47306 ‘The Sapper’

There is light at the end of the tunnel as we are again putting things back on 47306 rather than taking them off! This last couple of months has seen a lot of effort going into the Radiator fan grills which open and close slats by hydraulic oil pressure. All four frames have been lowered down off the top of the loco. The first two B side have been stripped down, cleaned back to bare metal, primed and painted, including the many slats. Reassembled and manually tested prior to being installed back over the top of the radiators. The second set of frames and slats ‘A side’ are already progressing and receiving the same treatment.

With the generator out, the whole area underneath has been cleaned, painted and new rubber insulation laid around the cable junction area. Further floor panels have been cut, welded and painted around the engine area completing the floor repairs in the engine room.

Work also continues on repairing exhaust system parts as well as making the new engine room roof strips that the roof sits on and seals.

Cab1 work continues on the wetter days and the head code panel has also now been removed from this end too which like Cab 2 is badly corroded so will probably be replaced.

The news from the contractors that took receipt of the generator set, is outside of the banding starting to come away, the generator is in relatively good condition. It will still undergo a cleaning process, minor repairs, insulation varnishing, baking before being reassembled and returned.

There is of course lots of sanding, filling and painting along with a host of other minor repair work going on. Volunteer man hours of work on 47306 now exceed 3,500 since start of overhaul Nov 2016!













  • Comments(0)//blog.bodmin-heritage-diesels.co.uk/#post27